Sorin Inc., a company that produces and sells a single product, has provided its contribution format income statement for January Sales (3,500 units) Variable expenses Contribution margin Fixed expenses Net operating income $94,500 41,580 52,920 40,700 $12,220 If the company sells 3,800 units, its total contribution margin should be closest to: (Do not round intermediate calculations.) Multiple Choice $52,920 $57456 $71,700 $13,267

Nocum Corporation has provided the following contribution format income statement. Assume that the following information is within the relevant range. Sales (3,000 units) Variable expenses Contribution margin Fixed expenses Net operating income $120,000 90,000 30,000 21,000 $ 9,000 If sales decline to 2,900 units, the net operating income would be closest to Multiple Choice $29,000 $1,000 $8,700 $8,000

Carver Corporation produces a product which sells for $40. Variable manufacturing costs are $18 per unit. Fixed manufacturing costs are $5 per unit based on the current level of activity, and foxed sellig and administrative costs are $4 per unit. A selling commission of 15% of the selling price is pad on each unit sod.The contitution margin per unit is Multiple Choice $7 $17 O $22 $16
Answer #1

1) Total Contribution margin on 3800 Units = 52920/3500*3800 = $57456

So answer is b) $57456

2) Net operating income on 2900 Units :

Total Contribution margin on 2900 Units = 30000/3000*2900 = 29000

Net operating income = 29000-21000 = $8000

So answer is d) $8000

3) Contribution margin per unit = Sale per unit-Variable cost per unit

= 40-(40*15%+18)

Contribution margin per unit = $16

So answer is d) $16